# Revisiting 360 Depth Estimation with PanoGabor: A New Fusion Perspective

### v4

@@ IV-B Comparison Analysis
IV-C Ablation Study
To demonstrate the effectiveness of the main proposed components, we follow Unifuse [2] and Elite360D [7] to design the ablation study on the Matterport3D dataset [13]. All the models have the same conditions and are trained to converge. As illustrated in Fig. 2, we remove the fusion stage from the full framework to serve as the baseline for conducting the ablation study. The experimental results of the baseline are presented as ‘a’ in Table LABEL:tab:ablation. We further add the proposed CS-UFM module, a version without the PanoGabor layer embedded as ‘b’. Index ‘c’ refers to the use of the full CS-UFM module. Index ‘d’ indicates that we employ the proposed spherical gradient constraint to stabilize the orientation sensitivity of the Gabor filters. Additionally, in ‘e’, all PanoGabor layers are replaced with normal Gabor filters to further verify the effectiveness of the designed PanoGarbor.
@@ IV-C Ablation Study
Effectiveness of the PanoGabor layer. Comparing the b and e in Table LABEL:tab:ablation and Fig. 10, we can observe that benefiting from the frequency domain transformation along with the capability of reducing surface texture interference, the normal Gabor layer can benefit the fusion process. From d and e, we can observe that our designed PanoGabor performs better than the normal ones due to the distortion-aware designs.
@@
Effectiveness of the spherical gradient constraint. Recently, some works [4, 7] prefer to employ a gradient-based constraint to sharpen the depth edges. This strategy aims to extract the horizontal/vertical structures to design the objective function. However, since the Gabor filters are orientation-sensitive, we design the spherical gradient constraint to incorporate the panoramic geometry to stabilize this orientation sensitivity. From Table LABEL:tab:ablation and Fig. 10, we can observe that the PanoGabor with the proposed spherical gradient constraint performs much better than that without.
Various Backbones. Following Elite360D [7], we also adopt various backbones for our framework. From Table IV, we can observe that our scheme can benefit from employing more powerful backbones.
Trade-off between performance and efficiency. From Table V, Elite360D has obtained a significant improvement by employing EfficientNet-B5 as a feature extractor. However, it brings unbearable memory costs. In contrast, our approach can achieve superior performance just by replacing a group of input features without additional memory cost. This also demonstrates the robustness and scalability of our PGFuse.
TABLE IV: Ablation study on various backbones.
@@
[b] Datasets Backbone Abs Rel $\downarrow$ Sq Rel $\downarrow$ $\mathrm{RMSE}\downarrow$ $\delta_{1}(\%)\uparrow$ ResNet-18 0.1125 0.0977 0.4933 87.78 M3D ResNet-34 0.1067 0.0858 0.4624 89.06 ResNet-50 0.1058 0.0841 0.4618 88.93
@@
TABLE V: More advanced performance comparison results. * represent that Elite360D employs Efficient-B5 as backbone. indicates that we replace one of the cubemap features with extracted ERP features.
@@ V Conclusion
In this paper, we have proposed a Gabor-based framework, termed PGFuse for 360 depth estimation. Since the existing fuse-based scheme ignores the reintroduced distortion when integrating other representations into ERP format, our proposed PGFuse embeds a Gabor-based layer into a unidirectional fusion module to explicitly deal with the reintroduced distortion. Specifically, to make the Gabor filters adapt to 360 images, we first introduce a latitude-aware representation to model the panoramic distortion distribution pattern. Then we leverage this representation to generate customized Gabor filters. These tuned PanoGabor can enhance the depth cues (i.e., resistant to surface texture and lighting interference, while preserving structural details) in the frequency domain and achieve effective distortion awareness. Therefore, our fusion module can deliver spatial-continuous and content-consistent ERP features. Extensive experiment results demonstrate that our approach outperforms current SoTA approaches.

### v5

@@ IV-C Ablation Study
Effectiveness of the PanoGabor. Comparing b and e in Table LABEL:tab:ablation and Fig. 9, we can observe that, benefiting from the frequency domain transformation along with the capability of reducing surface texture interference, the normal Gabor layer can benefit the fusion process. From d and e, we can observe that our designed PanoGabor performs better than the normal ones due to the distortion-aware designs.
Figure 11: Qualitative comparison of feature responses with and without the Spherical Gradient constraint.
@@
Effectiveness of the spherical gradient constraint. Recently, some works [28, 3] prefer to employ a gradient-based constraint to sharpen the depth edges. This strategy aims to extract the horizontal/vertical structures to design the objective function. However, since the Gabor filters are orientation-sensitive, we design the spherical gradient constraint to incorporate the panoramic geometry to stabilize this orientation sensitivity.
From Table LABEL:tab:ablation, we observe that the PanoGabor with the proposed spherical gradient constraint outperforms the version without it (i.e., settings (c) vs. (d) in Table LABEL:tab:ablation). Removing the constraint consistently degrades all evaluation metrics (e.g., RMSE rises from 0.4624 to 0.4767, $\delta_{1}$ drops from 89.06 to 88.51). Importantly, this degradation is consistent across both error- and accuracy-based measures, suggesting that orientation sensitivity introduces instability at both global and local levels.
As shown in Fig. 11, the feature responses without the constraint exhibit pronounced orientation-dependent variations: edges aligned with certain directions are overly amplified, while others are suppressed or distorted, resulting in visible artifacts such as uneven contrast and directionally biased activations. In contrast, applying the constraint produces smoother and more homogeneous responses across orientations, where structural boundaries are preserved more uniformly and spurious artifacts are greatly reduced. This visual evidence reveals the underlying cause of the observed performance drop and demonstrates that the spherical gradient constraint effectively enforces orientation-robust feature representations.
Figure 12: Qualitative comparison results on indoor (top) and outdoor (bottom) scenes.
Generalization Analysis. We evaluate the generalization ability of PGFuse on the SUN360 dataset [35], which contains diverse panoramic images without depth annotations. Since ground truth is unavailable, we report qualitative results in Fig. 12.
As shown in the outdoor examples, PGFuse yields plausible estimates of near-range structures such as buildings and ground surfaces, while predictions for distant regions are less precise. This discrepancy arises from the inherent differences between indoor and outdoor data: (1) indoor training datasets constrain the depth range to within 10 m, whereas outdoor scenes span substantially larger ranges; and (2) outdoor panoramas often include large homogeneous regions (e.g., sky) and more diverse illumination conditions. This illustrates the intrinsic challenge of transferring a model trained on indoor data (with depth limited to 10 m) to outdoor environments.
As shown in Fig. 12, PGFuse provides consistent and structurally reliable depth maps, with well-preserved object boundaries and relative depth relations. In comparison, Elite360D[3] also produces reasonable results, but tends to lose details in complex regions (highlighted in red boxes), particularly under outdoor settings. These findings confirm that PGFuse generalizes robustly within the indoor domain, while still delivering plausible outdoor predictions despite the increased difficulty.
TABLE IV: Sensitivity of model performance to Gabor parameter initialization.
@@
Init. Gabor Param. $\mathrm{RMSE}\downarrow$ Sq Rel $\downarrow$ Abs Rel $\downarrow$ $\delta_{1}(\%)\uparrow$
Random Non-convergence
Xavier(U) 0.4856 0.0933 0.1127 87.34
Gabor Init. 0.4848 0.0917 0.1089 88.25
Ours 0.4624 0.0858 0.1067 89.06
@@
Analysis of initialization sensitivity for Gabor parameters. In Table LABEL:tab:sens, we evaluate four initialization strategies for Gabor parameters: random initialization leads to non-convergence; Xavier initialization [14] converges but gives suboptimal accuracy since it ignores the structured nature of Gabor filters; handcrafted Gabor initialization [4] provides frequency-selective priors but neglects panoramic distortions; and our distortion-aware initialization achieves the best performance by aligning orientations and frequencies with spherical geometry. These results confirm that PGFuse is sensitive to initialization, but the proposed distortion-aware scheme can effectively alleviate this sensitivity, stabilize training, and improve accuracy.
TABLE V: Ablation study on various backbones.
Dataset Backbone Abs Rel $\downarrow$ Sq Rel $\downarrow$ $\mathrm{RMSE}\downarrow$ $\delta_{1}(\%)\uparrow$
M3D ResNet-18 0.1125 0.0977 0.4933 87.78
ResNet-34 0.1067 0.0858 0.4624 89.06
ResNet-50 0.1058 0.0841 0.4618 88.93
  • The bold values indicate the best performance.
Various Backbones. Following Elite360D [3], we also adopt different backbones for our framework. From Table V, we can observe that our scheme can benefit from employing more powerful networks. In particular, PGFuse demonstrates strong scalability and can effectively leverage stronger backbones. Unlike UniFuse, which suffers from limited performance gains due to reintroduced distortion when upgrading the backbone[17], PGFuse maintains consistent improvements across ResNet-18, ResNet-34, and ResNet-50, clearly validating the robustness of our design against structural distortion.
TABLE VI: More advanced performance comparison results. * represent that Elite360D employs Efficient-B5 as backbone. indicates that we replace one of the cubemap features with extracted ERP features.
@@
Trade-off between performance and efficiency. From Table VI, Elite360D has obtained a significant improvement by employing EfficientNet-B5 as a feature extractor. However, it brings unbearable memory costs. In contrast, our approach can achieve superior performance just by replacing a group of input features without additional memory cost. This also demonstrates the robustness and scalability of our PGFuse.
Discussion of Projection Trade-offs. Although PGFuse is formulated as a two-stage distortion-optimization pipeline to address different distortion problems in complementary domains, the same pipeline can be instantiated with alternative projection choices (e.g., ERP-only or ERP+cubemap), providing practical flexibility to balance global consistency, local detail accuracy, and computational cost according to specific application requirements.
Specifically, we evaluate three settings: ERP-only, dual-cubemap (our default configuration), and ERP+cubemap. The ERP-only setting estimates panoramic depth directly in the full ERP domain. It provides a globally consistent representation and a lower computational cost. However, PanoGabor would have to be incorporated into the backbone to explicitly handle panoramic distortion. This modification changes the architecture on which the weights were pretrained and typically weakens the effect of pretraining; thus, this setting struggles to capture local details. In contrast, the dual-cubemap setting adopts a two-stage distortion-optimization pipeline. It first extracts features in the low-distortion cubemap domain and then refines them in the ERP domain with PanoGabor. This design largely reduces the negative effects of panoramic distortion while keeping the computational cost acceptable. It produces finer local details and comparable global consistency. The ERP+cubemap setting combines the strengths of the two settings above. However, it requires two separate backbones to extract features from the ERP and cubemap representations, which leads to significantly higher computational cost.
@@ V Conclusion
In this paper, we have proposed a Gabor-based framework, termed PGFuse, for 360 depth estimation. Since the existing fuse-based scheme ignores the reintroduced distortion when integrating other representations into ERP format, our proposed PGFuse embeds a Gabor-based layer into a unidirectional fusion module to explicitly deal with the reintroduced distortion. Specifically, to make the Gabor filters adapt to 360 images, we first introduce a latitude-aware representation to model the panoramic distortion distribution pattern. Then we leverage this representation to generate customized Gabor filters. These tuned PanoGabor can enhance the depth cues (i.e., resistant to surface texture and lighting interference, while preserving structural details) in the frequency domain and achieve effective distortion awareness. Therefore, our fusion module can deliver spatial-continuous and content-consistent ERP features. Extensive experiment results demonstrate that our approach outperforms current SoTA approaches.